January 13, 1998

PHY 742 – Quantum Mechanics II

General Outline

1. Magnetic effects — L&L Chapter XV 2. Atomic structure

(a) Hartree-Fock theory

(b) Density functional theory

3. Molecules

4. Dirac equation